OpenStudy (anonymous):

Is 11101.111 (base2) EF in base 16?

OpenStudy (cwrw238):

hmmmmm 11101 = 2^0 + 0 + 4 + 8 + 16 = 29 in base 10 = 1D in base 16 - if my memory serves me right

ganeshie8 (ganeshie8):

i think he means 11101111

OpenStudy (cwrw238):

oh right

OpenStudy (anonymous):

The problem is written as follows : 11101.111 and I am asked to write the following base 2 numbers in base 16.

OpenStudy (phi):

to quickly change from base 2 to base 16, group 4 digits together 1 1101 . 1110 1 13 . 14 or using letters 1D.E

ganeshie8 (ganeshie8):

well, then its not EF for sure. to convert from base2 to base16 you can group binary digits by 4

OpenStudy (anonymous):

I was confused to the decimal point. I didn't know why it was there.

OpenStudy (cwrw238):

1110 = E 1111 = F

OpenStudy (anonymous):

So 100.1100101 would be 365?

OpenStudy (phi):

it is a fractional number base 2 See http://www.wolframalpha.com/input/?i=11101.1110+base+2+to+base+16

ganeshie8 (ganeshie8):

it wud be 4.CA i think, you need to group by 4

OpenStudy (anonymous):

I did group by 4.

OpenStudy (anonymous):

110.1100101 is the correct base 2 number. I apologize I may have typed it in incorrectly.

OpenStudy (phi):

100.1100101 would be 365? 0100 . 1100 1010 4 . C A

OpenStudy (phi):

then 6.ca

OpenStudy (anonymous):

HOW!

ganeshie8 (ganeshie8):

110.1100101 is same as 00000000000110.110010100000000000

OpenStudy (phi):

it helps if you know your powers of 2 (memorized to 2^10 !) 8 4 2 1 0 1 1 0

OpenStudy (phi):

the first digit of the fraction, from 1 1 0 0 is 8+4=12 or C 8 4 2 1

OpenStudy (anonymous):

So 6CA?

OpenStudy (phi):

you mean 6.ca with the "decimal" point

OpenStudy (anonymous):

>_< I guess? I'm sorry guys I am terrible at math.

OpenStudy (cwrw238):

i must remember the groups of four thing

OpenStudy (anonymous):

Why add leading zeros prior to the right most 1 in the problem?

OpenStudy (phi):

not required, but it reminds you of what the 4th digit is.

OpenStudy (anonymous):

It's the difference of if it's an A or not.

ganeshie8 (ganeshie8):

you dont have to if you dont wanto use the short cut. for using the short cut, you need four digits, so you need to add it

OpenStudy (phi):

it is more important for the fractional end: 0.1 must be written as 0.1000 so you write 0.8 rather than 0.1 (base 16)

OpenStudy (anonymous):

How can I get better at this stuff?

ganeshie8 (ganeshie8):

understand positional number system. once you grasp it you can do arithmetic in any numbersystem octal/binary/hexa... anything